# Peekaboo: Interactive Video Generation via Masked-Diffusion

### 12 Dec 2023 v1

@@ 5.2 Qualitative Results
(f) A ship docked at a port.
@@
6 Conclusion
Appendix A Implementation details
@@
Appendix B Ablation studies
B.1 Sensitivity to $t$
Figure 8: Sensitivity to frozen steps We plot AP50 against number of frozen steps $t$ for ModelScope and ZeroScope. The radius of the marker is proportional to the coverage. We find that increasing $t$ increases AP50 at the risk of losing coverage, i.e. degrading quality.
In fig 8, we present results on varying the $t$ parameter for generation on the IMC dataset. As $t$ increases, AP50 increases, but coverage decreases
B.2 More results on masking
In Tab 3, we present detailed results from Fig4.
@@ B.2 More results on masking
Appendix C More videos
We have uploaded the videos of the results presented in the main paper with our supplementary material. We also append more video results in the supplementary material for the reader.
Appendix D Dataset Curation and filtering
D.1 IMC
D.1.1 Prompts
List of prompts:
@@ D.1.1 Prompts
D.1.2 Generating the bounding Boxes
@@
Appendix E Limitations
@@
Appendix F Societal Impact

### 19 Apr 2024 v2

@@ Appendix A Text to image synthesis
(b) A ship docked at a port.
@@
Appendix B Implementation details
@@
Appendix C Ablation studies
C.1 Sensitivity to $t$
Figure 8: Sensitivity to frozen steps: We plot AP50 against number of frozen steps $t$ for ModelScope and ZeroScope. The radius of the marker is proportional to the coverage. We find that increasing $t$ increases AP50 at the risk of losing coverage, i.e. degrading quality.
In fig 8, we present results on varying the $t$ parameter for generation on the IMC dataset. As $t$ increases, AP50 increases, but coverage decreases. We hypothesize that this is because of selective attention masking for longer steps affecting the diffusion process. As the base models, were not trained with selective attention the model output quality degrades with larger number of steps.
C.2 More results on masking
We supplement the results of Fig 4 with a full results table in Tab 3. As we can see from AP50, Cross Attn mask is responsible for the majority of the control as it determines the location of the foreground object in the canvas. While Self Attn mask is responsible for maintaining generated video quality as observed by Coverage scores.
@@ C.2 More results on masking
Appendix D Effect of subjects on video generation
Figure 9: Effect of subjects on video generation quality: We plot mIoU on LaSOT against different subjects present in the dataset. We observe that Peekaboo either retains or improves generation quality for majority of subjects.
Evaluating quality across foreground subject depends heavily on the base model’s training data. However, as we show in Fig 6, using masked diffusion in Peekaboo can overcome model failures to some extent. We further evaluate the effect of subjects on controlled video generation by Peekaboo against LaSOT dataset. Fig 9 shows that Peekaboo overcomes the base model’s training bias and improves the control in video generation. Interestingly, Peekaboo performance is better for some subject classes than other, perhaps it is a factor of disentanglement of CLIP embeddings used in cross-attention. Studying the cause of this imbalance warrants further research and we encourage the community to do so.
Appendix E More videos
We have uploaded the videos of the results presented in the main paper with our supplementary material and are available on the project webpage. Please use VLC media player for viewing the videos. We also append more video results in the Fig 10 for the reader.
Appendix F Dataset Curation and filtering
F.1 IMC
In this section, we provide finer details of generating IMC dataset.
F.1.1 Prompts
We provide the list of prompts used in IMC dataset.
@@ F.1.1 Prompts
F.1.2 Generating the bounding Boxes
@@
Figure 10: More interactive video generation results
Appendix G Limitations
@@
Appendix H Societal Impact